OpenStudy (khalilforthewin):

Maria and Ming start at the same point and drive in opposite directions. Maria drives 34 mph and Ming drives 42 mph. How far apart will they be in 2 hours? A. 110 miles B. 162 miles C. 76 miles D. 152 miles I do not understand this please help

OpenStudy (asad786):

34 times 2 =

OpenStudy (anonymous):

wiz 68

OpenStudy (anonymous):

wiz khalifa

OpenStudy (asad786):

yes and now whats 42times 2

OpenStudy (asad786):

the answer is the last one

OpenStudy (khalilforthewin):

?

OpenStudy (khalilforthewin):

42times2 is 84

OpenStudy (khalilforthewin):

@Asad786

OpenStudy (khalilforthewin):

yea ik then you subtract but it isnt up there

OpenStudy (khalilforthewin):

@Amberpatterson @DaWizjr @k_lynn @MasterprojectHD

OpenStudy (khalilforthewin):

@israr000

OpenStudy (khalilforthewin):

@chosenmatt

OpenStudy (khalilforthewin):

@crystelle

OpenStudy (khalilforthewin):

@bib

OpenStudy (khalilforthewin):

@bibby

OpenStudy (khalilforthewin):

@JulieBeans

OpenStudy (khalilforthewin):

@tjb69812

OpenStudy (anonymous):

Its D.

OpenStudy (khalilforthewin):

how

OpenStudy (khalilforthewin):

@welshfella

OpenStudy (welshfella):

they are driver apart at (34+420 mph

OpenStudy (welshfella):

* 34 + 42 = 76 mph so you need to find how far they go in 2 hours using this speed

OpenStudy (khalilforthewin):

so its 34time2-42times2

OpenStudy (khalilforthewin):

but that doesnt work

OpenStudy (welshfella):

no - add the speeds then use distance = speed * time

OpenStudy (khalilforthewin):

ok so it does have something to do with science

OpenStudy (anonymous):

Its D -_-. Because since its (PER HOUR) you would muiltiply 34 and 42 by 2. Then add the sum.

OpenStudy (khalilforthewin):

i thought distance equal speed over time

OpenStudy (khalilforthewin):

@welshfella

OpenStudy (anonymous):

So 34*2 = 68 and 42*2 = 84. 68 + 84 = 152

OpenStudy (anonymous):

I believe.

OpenStudy (welshfella):

No the 2 speeds are 34 and 42

OpenStudy (khalilforthewin):

did u use the d=speed/time

OpenStudy (welshfella):

oh - i see - yes thats right

OpenStudy (khalilforthewin):

he is right?

OpenStudy (welshfella):

yes 152 is right and speed = distance / time distance = speed * time

OpenStudy (welshfella):

distance = 76 * 2 = 152

OpenStudy (khalilforthewin):

oh ok

OpenStudy (khalilforthewin):

thanks guys

OpenStudy (welshfella):

the 76 comes from adding the 2 speeds 34 and 42

OpenStudy (welshfella):

yw